Grade 12Mechanics

An apple gives 21 kilo joule energy to a boy how much height he can climb by using this energy if its efficiency is 28% mass of boy is 40 kg

Arti Bhagwan tijare
9 Years agoGrade 12

1 Answer

Lavina singh
9 Years ago
21 kJ energy=21000 J
energy utilised = 28/100 * 21000 = 5880 J
mgh=energy
mgh=5880J
40*10*h=5880
h=1.47 m